1. Suppose we are testing the difference between two proportions at the 0.05 level of significance. If the computed z is –1.07, what is our decision?
A) Reject the null hypothesis
B) Do not reject the null hypothesis
C) Take a larger sample
D) Reserve judgment

2. Suppose that an automobile manufacturer designed a radically new lightweight engine and wants to recommend the grade of gasoline that will have the best fuel economy. The four grades are: regular, below regular, premium, and super premium. The test car made three trial runs on the test track using each of the four grades and the miles per gallon recorded. At the 0.05 level, what is the critical value of F used to test the hypothesis that the miles per gallon for each fuel is the same.
Kilometers per liter
Regular Below Regular Premium Super Premium
39.31 36.69 38.99 40.04
39.87 40.00 40.02 39.89
39.87 41.01 39.99 39.93
A) 1.96
B) 4.07
C) 2.33
D) 12.00

3. Three different fertilizers were applied to a field of celery. In computing F, how many degrees of freedom are there in the numerator?
A) 0
B) 1
C) 2
D) 3

4. Suppose a package delivery company purchased 14 trucks at the same time. Five trucks were purchased from manufacturer A, four from B and five from manufacturer C. The cost of maintaining each truck was recorded. The company used ANOVA to test if the mean maintenance cost of the trucks from each manufacturer were equal. To apply the F test, how many degrees of freedom are in the denominator?
A) 2
B) 3
C) 11
D) 14


5. In ANOVA, an F statistic is used to test a null hypothesis such as:
A)
B)
C)
D)


6. An electronics company wants to compare the quality of their cell phones to the cell phones from three competitors. They sample 10 phones from each company and count the number of defects for each phone. If ANOVA were used to compare the average number of defects, the treatments would be defined as:
A) the number of cell phones sampled.
B) the average number of defects.
C) The total number of phones
D) The four companies.



7. The chi-square distribution can assume
A) only positive values.
B) only negative values.
C) negative and positive values or zero.
D) only zero.


8. A sample of 100 production workers is obtained. The workers are classified by gender (male, female) and by age (under 20, 20–29, 30–39 and 40 or over). How many degrees of freedom are there?
A) 0
B) 3
C) 6
D) 5


9. For any chi-square goodness-of-fit problem, the number of degrees of freedom is found by
A) n – k – 1.
B) k – 1.
C) n + 1.
D) n + k.


Answer by stanbon(57219)

1. Suppose we are testing the difference between two proportions at the 0.05 level of significance. If the computed z is –1.07, what is our decision?
The p-value would be P(-10 < z < -1.07) = 0.1423.. or more than 14%
You would not reject Ho because the p-value is > 5%.
----------------------
A) Reject the null hypothesis
B) Do not reject the null hypothesis
C) Take a larger sample
D) Reserve judgment
---------------------------------------

2. Suppose that an automobile manufacturer designed a radically new lightweight engine and wants to recommend the grade of gasoline that will have the best fuel economy.
The four grades are: regular, below regular, premium, and super premium.
The test car made three trial runs on the test track using each of the four grades and the miles per gallon recorded.
-----------------------------------------
At the 0.05 level, what is the critical value of F used to test the hypothesis that the miles per gallon for each fuel is the same.
That is a right-tail test with alpha=5%
----------------------------
Comment: I don't have an F-distribution table to check this answer.
----------------------------
Kilometers per liter
Regular Below Regular Premium Super Premium
39.31 36.69 38.99 40.04
39.87 40.00 40.02 39.89
39.87 41.01 39.99 39.93
A) 1.96
B) 4.07
C) 2.33
D) 12.00
---------------------
3. Three different fertilizers were applied to a field of celery. In computing F, how many degrees of freedom are there in the numerator?
df = k-1 where k is the number of samples.
Ans: 2
---------------------
A) 0
B) 1
C) 2
D) 3
-----------------------------
4. Suppose a package delivery company purchased 14 trucks at the same time. Five trucks were purchased from manufacturer A, four from B and five from manufacturer C. The cost of maintaining each truck was recorded. The company used ANOVA to test if the mean maintenance cost of the trucks from each manufacturer were equal. To apply the F test, how many degrees of freedom are in the denominator?
----------
Since the samples have a different number of elements, df = N-k = 14-3 =11
----------
Ans:
A) 2
B) 3
C) 11
D) 14
-------------------

5. In ANOVA, an F statistic is used to test a null hypothesis such as:
A) The means are all equal (three or more samples)
B) The proportions are all equal (three or more samples)
C)
D)
----------------------

6. An electronics company wants to compare the quality of their cell phones to the cell phones from three competitors. They sample 10 phones from each company and count the number of defects for each phone. If ANOVA were used to compare the average number of defects, the treatments would be defined as:
A) the number of cell phones sampled.
B) the average number of defects.
C) The total number of phones
D) The four companies.
I would say that none of these is correct. Treatments are the conditions
the objects being studied are randomly assigned to---for example some
subjects will receive a drug, others will receive a placebo. That
involves testing with two treatments.
None of those factors are involved in the example you are given.
------------------------------------


7. The chi-square distribution can assume
A) only positive values.
B) only negative values.
C) negative and positive values or zero.
D) only zero.
Ans: A because the data is # of observed occurrences.
--------------

8. A sample of 100 production workers is obtained. The workers are classified by gender (male, female) and by age (under 20, 20–29, 30–39 and 40 or over). How many degrees of freedom are there?
Ans: df=(2-1)(4-1) = 3
A) 0
B) 3
C) 6
D) 5
-------------------------

9. For any chi-square goodness-of-fit problem, the number of degrees of freedom is found by
Ans: k-1 where k is the number of data-pairs.
------------
A) n – k – 1.
B) k – 1.
C) n + 1.
D) n + k.

========================
Cheers,
Stan H.

--------------------------------------------------------------------------------

Textbook information: ISBN , Title Applied Statistics in Business and Economics , Chapter 13, Problem #13.32What's this?